OpenStudy (anonymous):

determine the formula for the Nth term of the following sequences, 2,3,6,11,18...

OpenStudy (solomonzelman):

\(\large\color{slate}{ 3=1^2+2 }\) \(\large\color{slate}{ 6=2^2+2 }\) \(\large\color{slate}{ 11=3^2+2 }\) \(\large\color{slate}{ 18=4^2+2 }\)

OpenStudy (solomonzelman):

converting this to a sequence notation, we get: \(\large\color{slate}{ a_2 =(2-1) ^2+a_1 }\) \(\large\color{slate}{ a_3 =(3-1) ^2+a_1 }\) \(\large\color{slate}{ a_4 =(4-1) ^2+a_1 }\) \(\large\color{slate}{ a_5 =(5-1) ^2+a_1 }\)

OpenStudy (solomonzelman):

\(\large\color{slate}{ a_n= }\) ..... ??

OpenStudy (anonymous):

an =1^2 ? @SolomonZelman